Publications

Les articles marqués d’une ^{≪} * ^{≫} peuvent être consultés à l’adresse

http://www.lamfa.u-picardie.fr/bouc

*[1] S. Bouc. – Homologie de certains ensembles ordonnés. C. R. Acad. Sc. Paris, t.299(2) : 49–52, 1984. Série I.

*[2] S. Bouc. – Modules de Möbius. C. R. Acad. Sc. Paris, t.299(1) : 9–12, 1984. Série I.

*[3] S. Bouc. – Projecteurs dans l’anneau de Burnside, projecteurs dans l’anneau de Green, et modules de Steinberg généralisés. J. of Algebra, 139(2) : 395–445, June 1991.

*[4] S. Bouc. – Exponentielle et modules de Steinberg. J. of Algebra, 150(1) : 118–157, August 1992.

[5] S. Bouc. – Homologie de certains ensembles de 2-sous-groupes des groupes symétriques. J. of Algebra, 150(1) : 158–186, August 1992.

*[6] S. Bouc. – Construction de foncteurs entre catégories de G-ensembles. J. of Algebra, 183(0239) : 737–825, 1996.

*[7] S. Bouc. – Foncteurs d’ensembles munis d’une double action. J. of Algebra, 183(0238) : 664–736, 1996.

*[8] S. Bouc. – Bimodules, trace généralisée, et transferts en homologie de Hochschild. Preprint, 1997.

[9] S. Bouc. – Green-functors and G-sets, volume 1671 of Lecture Notes in Mathematics. Springer, 1997.

*[10] S. Bouc. – Résolutions de foncteurs de Mackey. In Group representations : Cohomology, Group Actions and Topology, volume 63, pages 31–83. AMS Proceedings of Symposia in Pure Mathematics, 1998.

*[11] S. Bouc. – Sur le centre des algèbres de groupe en caractéristique p. Expositiones Math., 4 : 375-384, 1999.

*[12] S. Bouc. – Le complexe de chaînes d’un G-complexe simplicial acyclique. J. of Algebra, 220 : 415–436, 1999.

*[13] S. Bouc. – Burnside rings. In Handbook of Algebra, volume 2, chapter 6D, pages 739–804. Elsevier, 2000.

*[14] S. Bouc. – Non-additive exact functors and tensor induction for Mackey functors, volume 144 of Memoirs. A.M.S., 2000.

*[15] S. Bouc. – Polynomial ideals and classes of finite groups. J. of Algebra, 229 : 153–174, 2000.

*[16] S. Bouc. – Tensor induction of relative syzygies. J. reine angew. Math., 523 : 113–171, 2000.

*[17] S. Bouc. – A remark on a theorem of Ritter and Segal. J. of Group Theory, 4 : 11–18, 2001.

*[18] S. Bouc and J. Thévenaz. – The group of endo-permutation modules. Invent. Math., 139 : 275–349, 2000.

*[19] S. Bouc. – The p-blocks of the Mackey algebra. Algebras and Representation Theory 6 : 515-543, 2003 .

*[20] S. Bouc. – Green functors, crossed G-monoids, and Hochschild constructions. Algebra Montpellier Announcements Issue 01-2002.

http://www.emis.de/journals/AMA/AMA01-2002/index01-2002.html

*[21] S. Bouc. – Hochschild constructions for Green functors. Comm. in Algebra, 31(1) : 419-453, 2003

*[22] S. Bouc. – The functor of rational representations for p-groups. – Adv. in Maths., 186 : 267-306, 2004

*[23] S. Bouc. – A remark on the Dade group and the Burnside group. – J. of Algebra, 279(1) : 180-190, 2004.

*[24] S. Bouc. – The Burnside dimension of projective Mackey functors. – Proceedings of the Symposium “Algebraic Combinatorics” : 107-120, ISSN 1880-2818, RIMS, Kyoto, 2004

*[25] S. Bouc. – Biset functors and genetic sections for p-groups. – J. of Algebra, 284 : 179-202, 2005.

*[26] S. Bouc. – Mini-cours au CIRM : Green functors. –

http://www.math.jussieu.fr/∼bouc

*[27] S. Bouc and N. Mazza. – The Dade group of (almost) extraspecial p-groups. – J. of Pure and Applied Algebra. 192 : 21-51, 2004.

*[28] S. Bouc. – The Dade group of a p-group. – Inv. Math., 164 : 189-231, 2006.

*[29] S. Bouc. – The functor of units of Burnside rings for p-groups. – Comm. Math. Helv., 82 : 583-615, 2007.

*[30] S. Bouc and E. Yalçin. – Borel-Smith functions and the Dade group. – J. of Algebra, 311 : 821-839, 2007.

*[31] S. Bouc and J. Thévenaz. – Gluing torsion endo-permutation modules.

– J. London Math. Soc., 78 n˚ 2 : 477-501, 2008.

*[32] S. Bouc and J. Thévenaz. – A sectional characterization of the Dade group. – J. of Group Theory, 11 n˚ 2 : 155-183, 2008.

*[33] S. Bouc. – Rational p-biset functors. J. of Algebra, 319 : 1776-1800, 2008.

*[34] S. Bouc and J. Thévenaz. – The poset of elementary abelian subgroups of rank at least 2. L’Enseignement Mathématique, 40 : 41-45, 2008.

*[35] S. Bouc. – Bisets as categories, and tensor product of induced bimodules. Applied Categorical Structures, 18, n˚ 5 : 517-521, 2010.

*[36] S. Bouc. – Gluing endo-permutation modules. J. of Group Theory, 12, n˚ 5 : 651-678, 2009.

*[37] S. Bouc and K. Rökaeus. – A note on the λ-structure on the Burnside ring Journal of Pure and Applied Algebra, 213 : 1316-1319, 2009.

*[38] S. Bouc. – Complexity and cohomology of cohomological Mackey functors. Adv. in Maths, 221 : 983-1045, 2009.

*[39] S. Bouc and J. Thévenaz. – The primitive idempotents of the p-permu-tation ring. J. of Algebra, 323 : 2905-2915, 2010.

*[40] S. Bouc. – On the Cartan matrix of Mackey algebras. Trans. AMS, 363, n˚ 8 : 4383-4399, 2011.

[41] S. Bouc. – Biset functors for finite groups, volume 1990 of Lecture Notes in Mathematics, Springer, 2010.

*[42] S. Bouc and J. Thévenaz. – Stabilizing bisets. Adv. in Maths, 229 : 1610-1639, 2012.

*[43] S. Bouc and R. Stancu. – The extension algebra of some cohomological Mackey functors. Adv. in Maths., 283 (2015) 51-87.

*[44] S. Bouc. – The slice Burnside ring and the section Burnside ring of a finite group. Compositio Math., 148 : 868-906, 2012.

*[45] S. Bouc. – The Roquette category of finite p-groups. J.E.M.S., 17 (2015) 2843-2886

*[46] S. Bouc. – A conjecture on B-groups. Math. Zeitschrift, 274 Issue 1-2 : 367-372, 2013.

*[47] S. Bouc, R. Stancu, and J. Thévenaz. – Simple biset functors and double Burnside ring. J.P.A.A., 217 : 546-566, 2013.

*[48] S. Bouc, R. Stancu, and J. Thévenaz. – Vanishing evaluations of simple functors. J.P.A.A. 218 : 218-227, 2014.

*[49] S. Bouc. – Fused Mackey functors. Geometriae Dedicata, 176 : 225-240 (2015).

[50] S. Bouc. – The poset of posets. Preprint, 2013. arXiv :1311.2219

*[51] S. Bouc. – Fast decomposition of p-groups in the Roquette category, for p > 2. RIMS Kôkyûroku (ISSN 1880-2818), 1872 : 113-121, 2014.

*[52] S. Bouc and J. Thévenaz. – The algebra of essential relations on a finite set. J. reine angew. Math., 712 (2016) 225-250

*[53] S. Bouc and A. Zimmermann. – On a question of Rickard on tensor product of stably equivalent algebras. Experimental Mathematics, 2017, vol. 26, n˚ 1, 31-44.

*[54] S. Bouc, R. Stancu, and P. Webb. – On the projective dimensions of Mackey functors. Algebras and Representation Theory (2017) 20 :1467-1481.

*[55] S. Bouc and J. Thévenaz. – The representation theory of finite sets and correspondences. Preprint, 2015. arXiv :1510.03034.

*[56] S. Bouc and J. Thévenaz. – The monoid algebra of all relations on a finite set. Preprint, 2015. arXiv :1511.05048.

*[57] S. Bouc. – Correspondence functors. RIMS Kôkyûroku (ISSN 1880-2818), 1967 : 59-67, 2015.

*[58] S. Bouc. – Extensions of simple cohomological Mackey functors. RIMS Kôkyûroku (ISSN 1880-2818), 1967 : 115-122, 2015.

*[59] S. Bouc and N. Romero. – The Whitehead group of (almost) extraspecial p-groups with p odd, J.P.A.A. 223 (2019) 86-107.

*[60] S. Bouc. – K-theory, genotypes, and biset functors, arXiv :1604.07703. Boletín de la Sociedad Matemática Mexicana 26 : 57-74 (2020).

*[61] S. Bouc. – Idempotents of double Burnside algebras, L-enriched bisets, and decomposition of p-biset functors. J. Alg. 504 (2018) 129-175.

*[62] S. Bouc and J. Thévenaz. – Correspondence functors and finiteness conditions. J. Alg. 495 (2018) 150-198.

*[63] S. Bouc and J. Thévenaz. – Correspondence functors and lattices. J. Alg. 518 (2019) 453-518.

*[64] S. Bouc and N. Romero. – The center of a Green biset functor. Pacific Journal of Maths 303 (2019) 459-490.

*[65] S. Bouc and J. Thévenaz. – The algebra of Boolean matrices, correspondence functors, and simplicity. J. Comb. Alg. 4 (2020) 215-267.

*[66] S. Bouc and H. Mutlu. – Monomial G-posets and their Lefschetz invariants. J. Alg. 527 (2019) 399436.

*[67] S. Bouc and J. Thévenaz. – Tensor product of correspondence functors. J. Alg. 558 (2020) 146-175.

*[68] S. Bouc and J. Thévenaz. – Simple and projective correspondence functors. Representation Theory Volume 25, Pages 224-264 (2021).

[69] S. Bouc and J. Thévenaz. – A central idempotent in the endomorphism algebra of a finite lattice. Preprint (2019). arXiv :1902.10818.

*[70] S. Bouc. – Relative B-groups. Doc. Math. 24, 2431-2462 (2019).

*[71] S. Bouc and Deniz Yılmaz. – Diagonal p-permutation functors. J. Alg. 556 (2020) 10361056.

*[72] S. Bouc. – Representations of finite sets and correspondences. Proceedings of the ICRA 2018. Contemp. Math. 758 (2020) 31-55.

*[73] S. Bouc. – A functorial presentation of units of Burnside rings. Results Math. (2021) 76 :58

*[74] S. Bouc and N. Romero. – Shifted functors of representation functors are semisimple. Math. Zeitschrift 300, 247-257 (2022)

*[75] S. Bouc and N. Romero. – Green fields. Expositiones Mathematicae 40 (2022) 351-356.

*[76] S. Bouc and J. Thévenaz. – Idempotents in the endomorphism algebra of a finite lattice. Order 39, 485-504 (2022).

*[77] S. Bouc. – Germs in a poset. Journal of Algebraic Combinatorics 55, 493-511 (2022).

*[78] S. Bouc. – Some simple biset functors. Journal of Group Theory 26 (2023) 1-27.

*[79] S. Bouc and J. Thévenaz. – Correspondence functors and duality. Journal of Algebra 617 (2023) 425-475.

*[80] S. Bouc and Deniz Yılmaz. Diagonal p-permutation functors, semisimplicity, and functorial equivalence of blocks. Adv. Maths 411 (2022) 108799.